What is the final salt concentration if 20kg of a solution containing 21 percent (m/m) salt is mixed with 10kg of a solution containing 12 percent (m/m) salt?
salt mass= 20kg x .21 kg/kg + 10kg x.12kg/kg= 4.2 kg + 1.2kg= 5.4 kg
total solution mass= 20 kg+ 10 kg= 30 kg
How do I calculate the final concentration?

Answer by Theo(13342)

you let x equal the concentration of salt in the final solution.
your formula becomes:

.21(20) + .12(10) = x(30)
this equation becomes:
4.2 + 1.2 = x(30)
this simplifies to:
5.6 = x(30)
divide both sides of this equation by 30 to get:
x = 5.6/30 = .1867 rounded to the fourth decimal digit.
confirm by multiplying 30 * .1867.
you should get something close to 5.6.
you'll get much closer or right on if you don't round.